# A semicircle law for the normalized Laplacian of sparse random graphs

### v1

@@ 2.1. The Erdős-Rényi graph $G(n,p_{n})$
A careful examination of the proof indicates that Theorem 2 holds for $np_{n}>10\log n$, although no attempt is made to optimize the numerical constant.
@@ 3. Preliminaries
For any symmetric metric $A,B$, we have
@@
where $d_{\text{BL}}$ is a metric for the weak convergence of measures [dudley2018real].
@@
$\leq\frac{1}{n}\operatorname{tr}((\mathbf{B}-\mathbf{A})^{2}).$
where the last inequality is from Lidskii’s theorem ([bai1999methodologies]).∎
@@ 4. Proof of Theorems 1 and 2
Assume that the matrices $\mathbf{\mathcal{L}}_{n}$ and $\widetilde{\mathbf{\mathcal{L}}}_{n}$ are defined as (3) and (8). Then, almost surely,
$\frac{1}{n}\operatorname{tr}(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n})^{2}-\frac{1}{n}\mathbb{E}(\operatorname{tr}(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n})^{2})=o(u_{n}^{-1})\qquad a.s.$
as $n\rightarrow\infty$.
@@ 4.1. Proof of Lemma 2
, where we use the pseudo-inverse convention of $\mathbf{D}_{n}^{-1}$.
@@ 4.2. Proof of Theorem 1
Recall the spectral limit of $\widetilde{\mathbf{\mathcal{L}}}_{n}$. By the result of Ding and Jiang [ding2010spectral], we have
@@ 4.3. Proof of Lemma 3
which depends only on the adjacency variables $\{X_{ij}\}$ of the random graph.
@@
$\Gamma:=E_{2}=\left\{d_{n,i}\in\left(\frac{u_{n}}{2},\frac{3u_{n}}{2}\right)\text{ for all }1\leq i\leq n\right\}.$
@@
When the event $E_{2}$ holds, changing one coordinate $(u,v)$ in $\{X_{ij}\}$ affects at most the degrees of vertices $u$ and $v$. Recalling that $d_{n,i}=\sum_{j\neq i}a_{ij}$, we have
@@
since $d_{n,i}\in[u_{n}/2,3u_{n}/2]$ under $E_{2}$. Hence, for $F:=n^{-1}f_{n}$, the typical Lipschitz constants satisfy
$c_{uv}\leq\frac{C_{1}}{n\,u_{n}^{2}}\qquad(u<v),$
whenever both configurations lie in $\Gamma$. Outside $\Gamma$, Obviously, when $d_{n,i}=O(1)$ for all $1\leq i\leq n$ then $\delta_{uv}$ change most. Therefore, recall that $F:=n^{-1}f_{n}$, we also have the crude worst-case bound
@@
for some absolute constant $C_{2}>0$ (only the order $O(n^{-1})$ matters).
@@
Using $\beta_{n}=u_{n}^{-2}$, we deduce from (20) that
@@
which establishes the desired concentration result.
@@ 5. Proof of Theorem 3
Since $d_{n,i}$ and $d_{n,j}$ are sums of independent Bernoulli random variables, by a Chernoff bound we have
@@
On $E_{3}^{c}$, we use the crude bound
@@
$\mathbb{E}\left[d_{BL}^{2}\left(\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{N}_{n}),\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{\widetilde{N}}_{n})\right)\right]\leq\bar{w}_{n}\cdot\frac{1}{n}\mathbb{E}\operatorname{tr}\left((\mathbf{N}_{n}-\mathbf{\widetilde{N}}_{n})^{2}\right)\leq O\left(\frac{1}{w_{n,\min}}\right)\to 0\quad\text{as n}\to\infty,$
@@
$\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{D}_{n}^{-1/2}\mathbf{A}_{n}\mathbf{D}_{n}^{-1/2})\Rightarrow\mu_{\text{sc}}\quad\text{in probability as }n\to\infty.$

### v2

@@ 2.1. The Erdős-Rényi graph $G(n,p_{n})$
A careful examination of the proof indicates that Theorem 2 holds for $C>36$, although no attempt is made to optimize the numerical constant.
@@ 3. Preliminaries
For any symmetric matrices $A,B$, we have
@@
$\leq\frac{1}{\sqrt{n}}\sqrt{\sum_{j=1}^{n}|\lambda_{j}(\mathbf{A})-\lambda_{j}(\mathbf{B})|^{2}}\leq\frac{1}{\sqrt{n}}\sqrt{\operatorname{tr}((\mathbf{B}-\mathbf{A})^{2})},$
where the last inequality is from Lidskii’s theorem (see, e.g., [bai1999methodologies, Lemma 2.3]). ∎
@@ 4. Proof of Theorems 1 and 2
Assume that the matrices $\mathbf{\mathcal{L}}_{n}$ and $\widetilde{\mathbf{\mathcal{L}}}_{n}$ are defined as (3) and (8) and $np_{n}\geq C\log n$. Then, asymptotically almost surely,
$\frac{1}{n}\operatorname{tr}(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n})^{2}-\frac{1}{n}\mathbb{E}(\operatorname{tr}(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n})^{2})=o(u_{n}^{-1}).$
@@ 4.1. Proof of Lemma 2
where we use the pseudo-inverse convention of $\mathbf{D}_{n}^{-1}$.
@@ 4.2. Proof of Theorem 1
By the result of Ding and Jiang [ding2010spectral], we have
@@ 4.3. Proof of Lemma 3
which depends only on the adjacency variables $\{X_{ij}\}$ of the random graph. Then
$F=\frac{1}{n}f_{n}=\frac{1}{n}\operatorname{tr}(\widetilde{\mathbf{\mathcal{L}}}_{n}-\mathbf{\mathcal{L}}_{n})^{2}.$
@@
$\Gamma:=\left\{d_{n,i}\in\left(\frac{u_{n}}{2},\frac{3u_{n}}{2}\right)\text{ for all }1\leq i\leq n\right\}.$
@@
When the event $\Gamma$ holds, changing one coordinate $(u,v)$ in $\{X_{ij}\}$ affects at most the degrees of vertices $u$ and $v$. Recalling that $d_{n,i}=\sum_{j\neq i}a_{ij}$, we have
@@
since $d_{n,i}\in[u_{n}/2,3u_{n}/2]$ under $\Gamma$. Hence, for $F:=n^{-1}f_{n}$, the typical Lipschitz constants satisfy
$c_{uv}\leq\frac{C_{1}}{n\,u_{n}^{2}},$
whenever the event $\Gamma$ holds.
Outside the typical event $\Gamma$, we distinguish two cases.
First, if $S_{u}=0$ or $S_{u}^{\prime}=0$, then necessarily $X_{uk}=0$ for all $k\neq u,v$. In this case the corresponding contribution vanishes identically, namely,
$\left|\sum_{k\neq u,v}^{n}\Big[\big(u_{n}^{-1}-(S_{u}S_{k})^{-1/2}\big)^{2}-\big(u_{n}^{-1}-(S_{u}^{\prime}S_{k})^{-1/2}\big)^{2}\Big]X_{uk}\right|=0.$
Otherwise, assume that $S_{u}S_{u}^{\prime}\neq 0$. Using the identities $x^{2}-y^{2}=(x+y)(x-y)$ and
$x^{-1/2}-y^{-1/2}=\frac{x-y}{xy\big(x^{-1/2}+y^{-1/2}\big)},$
we obtain
$\left|\sum_{k\neq u,v}^{n}\Big[\big(u_{n}^{-1}-(S_{u}S_{k})^{-1/2}\big)^{2}-\big(u_{n}^{-1}-(S_{u}^{\prime}S_{k})^{-1/2}\big)^{2}\Big]X_{uk}\right|$
$\leq$ $\left|\sum_{k\neq u,v}^{n}\Big(2u_{n}^{-1}-(S_{u}S_{k})^{-1/2}-(S_{u}^{\prime}S_{k})^{-1/2}\Big)\Big((S_{u}S_{k})^{-1/2}-(S_{u}^{\prime}S_{k})^{-1/2}\Big)X_{uk}\right|$
$\leq$ $c\left|\sum_{k\neq u,v}^{n}\Big((S_{u}S_{k})^{-1/2}-(S_{u}^{\prime}S_{k})^{-1/2}\Big)X_{uk}\right|$
$\leq$ $c\,\max_{k}S_{k}^{-1/2}\frac{|S_{u}-S_{u}^{\prime}|}{S_{u}S_{u}^{\prime}\big(S_{u}^{-1/2}+S_{u}^{\prime-1/2}\big)}\sum_{k\neq u,v}^{n}X_{uk}$
$\leq$ $c\,\max_{k}S_{k}^{-1/2}\frac{S_{u}}{S_{u}S_{u}^{\prime}\big(S_{u}^{-1/2}+S_{u}^{\prime-1/2}\big)}$
$=$ $c\,\max_{k}S_{k}^{-1/2}\frac{1}{S_{u}^{\prime}\big(S_{u}^{-1/2}+S_{u}^{\prime-1/2}\big)}.$
Under the pseudo-inverse convention, if there exists $1\leq i\leq n$ such that $d_{n,i}=O(1)$, then the increment $\delta_{uv}$ may attain its maximal size, and we obtain the crude bound
$\delta_{uv}=O(1).$
Therefore, recall that $F:=n^{-1}f_{n}$, we also have the crude worst-case bound
@@
for some absolute constant $C_{2}>0$.
@@
This also implies $\mathbf{D}_{n}$ is invertible asymptotically almost surely. Using $\beta_{n}=u_{n}^{-2}$, we deduce from (20) that
@@
which establishes the desired concentration result. ∎
@@ 5. Proof of Theorem 3
Since $d_{n,i}$ and $d_{n,j}$ are sums of independent Bernoulli random variables, by a Chernoff bound, we have
@@
On $E_{3}^{c}$, we use the crude bound and the pseudo-inverse convention to have
@@
$\mathbb{E}\left[d_{BL}^{2}\left(\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{N}_{n}),\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{\widetilde{N}}_{n})\right)\right]\leq\bar{w}_{n}\cdot\frac{1}{n}\mathbb{E}\operatorname{tr}\left((\mathbf{N}_{n}-\mathbf{\widetilde{N}}_{n})^{2}\right)=O\left(\frac{1}{w_{n,\min}}\right)\to 0\quad\text{as n}\to\infty,$
@@
$\hat{\mu}(\sqrt{\bar{w}_{n}}\mathbf{D}_{n}^{-1/2}\mathbf{A}_{n}\mathbf{D}_{n}^{-1/2})\Rightarrow\mu_{\text{sc}}\quad\text{in probability as }n\to\infty$